Technical Observing - LBC IQ Tests 20140318 UT

Observers: JHill, OKuhn, MEdwards (Tucson - remote room)
Telescope Operator: DHuerta (LBTO)

Summary

Wind was high for most of the night.

We qualified the new version of the LBC software and IA software on-sky, so we will keep it in place.

We did not get any data for LBC image quality issues.

John added a pointer to the telescope target to the allsky image in LBTplot.py.

Details

02:00 Closed early due to high winds

Open

06:55 Open in skies with lots of cirrus and moon

07:01 LBC Preset

07:05 DOFPIA, but blue side can't find pupils due to moon and cirrus.

07:25 Seeing is not so good, telescope is warm compared to ambient, lots of moon, lots of cirrus.......we are trying to get a long enough red observation to demonstrate that Z4 correction is working as expected.

07:35 Z4 corrections are being sent, and are being received by PSF, but the LBC GUI is not displaying the correction for long.

08:13 DOFPIA

08:17 lbcrangebal

Close

08:39 Close for increasing wind

11:27 Reconfigure for MODS1 spectroscopy in case the wind goes down

12:30 We were never able to reopen.

-- %USERSIG{JohnHill - 2014-03-18}%
Topic revision: r4 - 18 Mar 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback